You can't find this on the internet...
Fred is 36, and is twice as old as Bert was when Fred was as old as Bert is now. How old is Bert?


Answer by Theo(13342)

f = fred's age now.
b = bert's age now.
x = number of years

the problem statement is:

Fred is 36, and is twice as old as Bert was when Fred was as old as Bert is now.
How old is Bert?

fred as as old as bert is now x years ago.

the equation for that is f - x = b

fred is now twice as old as bert was then.

the equation for that is f = 2 * (b - x)

since b = f - x, then replace b in that equation with f - x to get:

f = 2 * (f - x - x)

combine like terms to get:

f = 2 * f - 2x)

simplify to get:

f = 2f - 4x

subtract f from both sides of that equation and add 4x to both sides of that equation to get:

4x = 2f - f

solve for f to get:

f = 4x

since f = 36, then 4x = 36.

solve for x to get x = 9

when x = 9, you get:

f = 36
b = 36 - 9 = 27
x = 9

now that you know these values, look at the problem statement again.

that statement is:

Fred is 36, and is twice as old as Bert was when Fred was as old as Bert is now.
How old is Bert?

fred is 36 and bert is 27.
9 years ago, fred was 27 and bert was 18.

9 years ago, fred was as old as bert is now.
9 years ago fred was 27.
bert is 27 now.
the statement is true.

fred is now twice as old as bert was then.
fred is 36 years old now.
bert was 18 years old 9 years ago.
36/18 = 2
the statement is true.

your solution is that bert is 27 years old now.